OpenStudy (anonymous):

The price of a technology stock was $9.72 yesterday. Today, the price fell to $9.63. Find the percentage decrease. Round your answer to the nearest tenth of a percent.

OpenStudy (anonymous):

it came out as a long number

OpenStudy (anonymous):

1.0093457943925233

OpenStudy (dangerousjesse):

\(9.63/9.72=x\) and \(100-x=?\)

OpenStudy (anonymous):

that came out a long number too... 0.9907407407407408

OpenStudy (dangerousjesse):

100-99.07407407407408=?

OpenStudy (anonymous):

0.9259259259259238

OpenStudy (anonymous):

The price of a cup of coffee has dropped to $2.20 today. Yesterday's price was $2.35. Find the percentage decrease. Round your answer to the nearest tenth of a percent. @DangerousJesse

OpenStudy (anonymous):

??